OpenStudy (anonymous):

When writing linear equations, how do you determine which form of a line to use?

OpenStudy (anonymous):

Depending on the information your given. If given two points, use Point-Slope form. If given an equation comparing x and x, use Slope-Intercept form If Given a point and an intercept, use Point-Slope form.

OpenStudy (andrewkaiser333):

what is the problem you are having to solve so i can help or is this the question needing answered

OpenStudy (anonymous):

If given an equation comparing x and y (sorry), use Slope-Intercept form

OpenStudy (anonymous):

The question was When writing linear equations, how do you determine which form of a line to use?

OpenStudy (anonymous):

I'm good at linear algebra I've realized haha

OpenStudy (anonymous):

lol kcool but im not

OpenStudy (anonymous):

I can teach you if you like all the basics, more advanced, and advanced aspects of linear algebra if you like. Right here! Although a 3 day a week, multiple example/problem solving, quiz, and testing form will be required and is the only way I teach and could take up to 3 weeks.

OpenStudy (andrewkaiser333):

oh thought it was a problem you did not understand i am sorry

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

no all i want to know is: When writing linear equations, how do you determine which form of a line to use?

OpenStudy (andrewkaiser333):

i am not good with explaining unless it is by giving and working out the problem step by step so you understand

OpenStudy (anonymous):

yess ok

OpenStudy (andrewkaiser333):

i am gonna leave i am not to good with that

OpenStudy (andrewkaiser333):

sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@Cosmichaotic how do you determine which form of a line to use? when you do lineaar equations

OpenStudy (anonymous):

Like I said, it totally depends on what information you are given. Like, lets say they give you an equation: 2x + 3y = 4 And that's all they give you. We want to rearrange this into Slope-Intercept Form: (y=mx+b) Which using algebra rules we get eventually too... y = -2/3x + 4/3 where we can say m = -2/3 and b = 4/3 m is your slope b is your y intercept That's why they call it slope-intercept form! After we plot the y-int at (0, 4/3) we can use rise/run with our slope to go down 2 and over to the right 3 units over and over to get our line.

OpenStudy (anonymous):

But let's say they give you 2 points. (0, 0) and (1, 1) and say graph the linear line that passes through these two points. You could use (0, 0) as (x, y) and (1, 1) as (x1, y1) and use point-slope form: y-y2 = m(x-x2) to determine the slope and then graph!

OpenStudy (anonymous):

Point-Slope Form, Slope-Incercept Form, and the Standard Form Ax + By + C = 0 can all be used.

OpenStudy (anonymous):

I hope this helps some.

OpenStudy (anonymous):

wait hold on k

OpenStudy (anonymous):

You should put: The form used to graph a linear line will depend on the type of information presented to determine the proper/most convenient method. Lol

OpenStudy (anonymous):

or should i put this : Linear equations are generally written in the standard form of the line Ax + By = C where A, B and C are real numbers and both of A and B are not equal to 0. To solve a set of linear equations it may be required to convert one of the equations to a form x = (C - By)/A or y = (C - Ax)/B which is then substituted for either x or y in the other equation. Also, it is quite easy to convert one form of the line to another with a minimal amount of calculation. A system of linear equations can be solved irrespective of which form of the line the equations are originally in.

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

That... is like... a freaking AMAZING answer @Queen_Bee1

OpenStudy (anonymous):

You just blew me away.

OpenStudy (anonymous):

Did you copy that from somewhere??

OpenStudy (anonymous):

aww thanx very much, nd lol yea i did copy and paste XD

OpenStudy (anonymous):

In the first part mention that C is a constant and can be any real number, including 0 Where {C|C = All real #'s} or something since you did so good explaining everything else haha. Way better than I could.